OpenStudy (anonymous):

Can someone please help me! How to write in simplest form? 1/x^2 * 1/x^3

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

$$\huge\frac{1}{x^2} * \frac{1}{x^3}$$

OpenStudy (skullpatrol):

$$\huge \frac{a}{b}*\frac{c}{d}=\frac{ac}{bd}$$

OpenStudy (skullpatrol):

$$\huge a^m*a^n=a^{m+n}$$

OpenStudy (jdoe0001):

keep in mind that \(\bf \Large \cfrac{1}{a^n}\implies a^{-n} \)

OpenStudy (asib1214):